A sample of 50 observations is taken from a normal population, with mu=100 and =10. If the population is finite with N=250. Find: a) 𝑃(𝑋̅ > 103) b) 𝑃(98 < 𝑋̅ < 101) c) Repeat above exercise with N=500


1
Expert's answer
2021-08-17T10:38:47-0400

Use the Finite Correction Factor for "n>0.05N"

a) "\mu=100, \sigma=10, N=250, n=50"


"\dfrac{\sigma}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}}=\dfrac{10}{\sqrt{50}}\sqrt{\dfrac{250-50}{250-1}}=\dfrac{20}{\sqrt{249}}"

"P(X>103)=1-P(Z\leq\dfrac{103-100}{\dfrac{20}{\sqrt{249}}})"

"\approx1-P(Z\leq2.367)\approx0.0090"

b) "\mu=100, \sigma=10, N=250, n=50"


"\dfrac{\sigma}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}}=\dfrac{10}{\sqrt{50}}\sqrt{\dfrac{250-50}{250-1}}=\dfrac{20}{\sqrt{249}}"


"P(98<X<101)=P(X<101)-P(X\leq98)"

"=P(Z<\dfrac{101-100}{\dfrac{20}{\sqrt{249}}})-P(Z\leq\dfrac{98-100}{\dfrac{20}{\sqrt{249}}})"

"\approx P(Z<0.7890)-P(Z\leq-1.5780)"

"\approx0.784944-0.057283\approx0.7277"


c) "\mu=100, \sigma=10, N=500, n=50"


"\dfrac{\sigma}{\sqrt{n}}\sqrt{\dfrac{N-n}{N-1}}=\dfrac{10}{\sqrt{50}}\sqrt{\dfrac{500-50}{500-1}}=\dfrac{30}{\sqrt{499}}"

"P(X>103)=1-P(Z\leq\dfrac{103-100}{\dfrac{30}{\sqrt{499}}})"

"\approx1-P(Z\leq2.2338)\approx0.0127"



"P(98<X<101)=P(X<101)-P(X\leq98)"

"=P(Z<\dfrac{101-100}{\dfrac{30}{\sqrt{499}}})-P(Z\leq\dfrac{98-100}{\dfrac{30}{\sqrt{499}}})"

"\approx P(Z<0.7446)-P(Z\leq-1.4892)"

"\approx0.771746-0.068215\approx0.7035"